Fermi National Laboratory
 

Operation Summary for January 19, 2011 - January 21, 2011

 

Wednesday Jan. 19

The day shift began with Operations monitoring the Tevatron (TEV) store 8429 (88.77E30) and the Recycler stash (341.8E10), with the Antiproton Source (Pbar) stacking (7.7E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI, MiniBooNE, and Meson
FTBF T-978 (CALICE) experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, Operations terminated store 8429 at 9:30 AM. Operations established store 8430 at 11:19 AM with an average initial luminosity (AIL) of 342.2E30. Unfortunately, store 8430 aborted when a low beta quadrupole magnet power supply (D0Q3) tripped off. Diagnosis of the power supply problem continued throughout the shift. A Pbar RF Anode power supply (ARF1) trip interrupted stacking for 30 minutes. The Neutrino and T-978 experiments ran all day.

During the evening shift, Operations prepared the D0 collision hall for access. EE Support personnel investigated and had D0Q3 repaired by 8:23 PM. Operations resumed ramping the TEV and began its checkout. Operations established store 8434 at 11:12 PM with an AIL of 259E30. Stacking, Neutrino, and T-978 beams continued as scheduled.

Thursday Jan. 20

The midnight shift began with Operations monitoring store 8434 (190.62E30) and the stash (77.0E10), with Pbar stacking (14.0E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations monitored the accelerator complex.

During the day shift, Operations terminated store 8434 at 2:09 PM. Operations established store 8435 at 3:32 PM with an AIL of 343.9E30. Operations reported that beam to NuMI and Pbar was interrupted for about 30 minutes due to a malfunctioning MI RF station. MiniBooNE took beam all shift. T-978 took beam until the MIRF trip and then elected to shut off for the rest of the day.

During the evening shift, Operations monitored the accelerator complex and delivered beam to the Neutrino users and Pbar.

Friday Jan. 21

The midnight shift began with Operations monitoring store 8435 (121.34E30) and the stash (243.0E10), with Pbar stacking (23.9E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations terminated store 8435 at 5:27 AM. Operations delivered beam to all users as per the schedule. Operations established store 8436 at 6:39 AM with an AIL of 342.73E30. Operations then established beam to T-978.

The Plans for Friday and the Weekend

The plans for today and the weekend are to stack, stash, and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 01/21/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory